# Aligning Model Properties via Conformal Risk Control

### v1

@@ Abstract
AI model alignment is crucial due to inadvertent biases in training data and the underspecified pipeline in modern machine learning, where numerous models with excellent test set metrics can be produced, yet they may not meet end-user requirements. Recent advances demonstrate that post-training model alignment via human feedback can address some of these challenges. However, these methods are often confined to settings (such as generative AI) where humans can interpret model outputs and provide feedback. In traditional non-generative settings, where model outputs are numerical values or classes, detecting misalignment through single-sample outputs is highly challenging.
In this paper we consider an alternative strategy. We propose interpreting model alignment through property testing, defining an aligned model $f$ as one belonging to a subset ${\mathcal{P}}$ of functions that exhibit specific desired behaviors. We focus on post-processing a pre-trained model $f$ to better align with ${\mathcal{P}}$ using conformal risk control. Specifically, we develop a general procedure for converting queries for a given property ${\mathcal{P}}$ to a collection of loss functions suitable for use in a conformal risk control algorithm. We prove a probabilistic guarantee that the resulting conformal interval around $f$ contains a function approximately satisfying ${\mathcal{P}}$.
Moreover, given the remarkable capabilities of modern AI models with large parameters and extensive training data, one might argue that alignment problems will naturally resolve, making alignment techniques unnecessary. However, we show that increasing the size of the training data or the number of parameters in a random feature model does not eliminate the need for alignment techniques when the pre-training data contains biased labels. We exhibit applications of our alignment methodology on a collection of supervised learning datasets for (shape-constrained) properties such as monotonicity and concavity. The general procedure is flexible and can be applied to a wide range of desired properties.
@@ 1 Introduction
The emergence of large foundation models has heightened the focus on the problem of alignment. Aligned models are artificial intelligences designed to pursue goals that align with human values, principles, and intentions (Leike et al., 2018; Ouyang et al., 2022; Hendrycks et al., 2023; Ngo et al., 2024). Although the alignment problem is predominantly examined in the context of potential artificial general intelligence (AGI), large language models (LLMs), and reinforcement learning (RL) agents, it also has roots in the modern machine learning pipeline (D’Amour et al., 2022). Motivated by this, we introduce a broader notion of alignment in this paper, extending beyond the aforementioned generative models to include even tabular regression models.
@@
As an example, consider a regression task where ${\mathcal{P}}$ represents models that are monotonically decreasing in a given feature. Constraining a prediction model during training to maintain monotonicity in this feature can be viewed as a form of alignment. However, ensuring that the predictions of a pre-trained model $f$ are monotonically decreasing in this feature can be more complex. This complexity arises particularly in non-generative settings where the user cannot update a pre-trained $f$ or obtain any outputs other than point predictions $f(X)$ for a given input $X$.
@@
In this work, we propose an approach to this problem of aligning a pre-trained model $f$ that is inspired by property testing (Ron, 2008; Goldreich, 2017) and utilizes conformal risk control (Angelopoulos et al., 2024). We propose a definition of functions that either satisfy a property ${\mathcal{P}}$ or are significantly far from satisfying ${\mathcal{P}}$, drawn from the field of property testing (Ron, 2008; Goldreich, 2017). Property testing focuses on designing testing algorithms for determining membership to ${\mathcal{P}}$ that are more efficient than the corresponding learning algorithms for ${\mathcal{P}}$ (Ron, 2008). From the perspective of modern deep learning models, we can imagine a user who wishes to determine whether a pre-trained model $f$ belongs to ${\mathcal{P}}$ but is unable to train a model of $f$’s size.
@@
Property testing algorithms generally perform local queries to determine, with high probability, whether a function has a given global property or is far from having that property. In this work, we consider mapping such queries for a given property of interest ${\mathcal{P}}$ to a corresponding set of loss functions. These loss functions can then be used in a conformal risk control procedure (Angelopoulos et al., 2024) to obtain a notion of alignment for ${\mathcal{P}}$. In particular, we prove that running this procedure results in a conformal interval around $f$ that contains a function close to ${\mathcal{P}}$.
@@
We demonstrate our methodology on real-world datasets for the properties of monotonicity and concavity. To obtain asymmetric conformal intervals around $f$ we provide a straight-forward generalization of the results of Angelopoulos et al. (2024) to multi-dimensional $\boldsymbol{\lambda}\in\mathbb{R}_{\geq 0}^{k}$. We focus the bulk of our experiments on monotonicity for the sake of exposition and because monotonicity constraints have been shown to be useful in promoting crucial aspects of alignment to human values, such as fairness and adherence to social norms (Wang and Gupta, 2020).
@@
Given the outstanding capabilities of modern AI models with substantially large numbers of parameters and training data, one may argue that the alignment problem may naturally disappear as such advances continue (Kaplan et al., 2020), rendering alignment techniques such as those presented in this paper unnecessary. However, another contribution of this paper is to refute this argument in a stylized setting, building on recent advances in the theory of linearized neural networks (Mei and Montanari, ; Misiakiewicz and Montanari, 2023). Specifically, we show that increasing the size of the training data or the number of parameters in a random feature model cannot help it satisfy a property ${\mathcal{P}}$, if the pre-training data has biased labels. Our simulations show that the result holds even if only a small fraction of the training labels are impacted by the bias.
@@
Summarizing our main contributions, we: (1) introduce an alignment perspective based on property testing, (2) use conformal risk control to post-process predictions of pre-trained models for better alignment, (3) generalize conformal risk control to multi-dimensional $\boldsymbol{\lambda}$, and (4) demonstrate that increasing training data and parameters in a random feature model does not eliminate the need for alignment. We discuss related work in Section 6, particularly our connections to Yadkori et al. (2024), who use conformal risk control to address LLM hallucinations (Ji et al., 2023).
@@ 2 Preliminaries
In this section we provide condensed overviews of property testing and conformal prediction and risk control. The main definitions and concepts we will use from property testing are $\varepsilon$-farness and proximity oblivious testers (POTs). We also discuss one-sided testers, but primarily to motivate these former two definitions; our proposed alignment procedure will mainly rely on the definition of proximity oblivious testers. We also discuss conformal prediction, mainly focusing on conformal risk control, as this will be the main tool used to achieve alignment in our setting.
2.1 Properties and property testing
Our perspective on alignment in this work is motivated by the field of property testing (Ron, 2008; Goldreich, 2017). The seminal work in this field by Blum et al. (1993) studied linearity testing of Boolean functions $f:\{0,1\}^{n}\to\{0,1\}$, and since then, there has been significant work on algorithms that are sublinear in $n$ for querying properties ${\mathcal{P}}$ such as low-degree polynomials (Kaufman and Ron, 2006; Bhattacharyya et al., 2009) and $k$-juntas (Blais, 2009). In this spirit, property testing algorithms generally make local queries to a small part of a large object, such as a function or graph. From these sublinear number of queries, the algorithm attempts to determine whether the object has the property at a global level or is significantly far from having the property.
Definition 1 ($\varepsilon$-farness).
For a function $f:{\mathcal{X}}\to{\mathcal{Y}}$, a distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, $\varepsilon>0$, and a property ${\mathcal{P}}$, we say $f$ is $\varepsilon$-far from ${\mathcal{P}}$ with respect to ${\mathcal{D}}$ if $\delta_{{\mathcal{P}},{\mathcal{D}}}(f)>\varepsilon$, where
@@ 2.1 Properties and property testing
That is, any function $g\in{\mathcal{P}}$ agrees with $f$ on inputs from ${\mathcal{D}}$ with probability at most $1-\varepsilon$ (Goldreich, 2017). While property testing often assumes ${\mathcal{D}}$ to be the uniform distribution, we will keep ${\mathcal{D}}$ general, referred to as the distribution-free setting (Halevy and Kushilevitz, 2007), as we are concerned with queries over arbitrary calibration data distributions.
Following the definition introduced by Halevy and Kushilevitz (2007), a distribution-free one-sided error tester for a property ${\mathcal{P}}$ is a probabilistic oracle machine $\mathsf{M}$, which is provided with a distance parameter $\varepsilon>0$, oracle access to an arbitrary function $f:\mathcal{X}\to\mathcal{Y}$, and oracle access to samples from a fixed but unknown distribution ${\mathcal{D}}$ over $\mathcal{X}$, and satisfies the following two conditions:
@@
    If $f$ satisfies $\mathcal{P}$, then $\Pr\{\mathsf{M}^{f,{\mathcal{D}}}=\text{Accept}\}=1$.
@@
    If $f$ is $\varepsilon$-far from $\mathcal{P}$ with respect to ${\mathcal{D}}$, then $\Pr\{\mathsf{M}^{f,{\mathcal{D}}}=\text{Accept}\}\leq\frac{1}{3}$.
2.1.1 Proximity oblivious testers
Given the definition of testers above, it is natural to expect that any reasonable tester must take in $\varepsilon$ as input. Many testing algorithms, however, rely on $\varepsilon$ only for the sake of determining the appropriate number of times some basic test should be invoked (Goldreich and Ron, 2008). The basic test itself does not use $\varepsilon$, so such basic tests are referred to as proximity oblivious (Goldreich and Ron, 2008).
Following Goldreich and Ron (2008), a proximity-oblivious tester (POT) for ${\mathcal{P}}$ is a probabilistic oracle machine $\mathsf{T}$ that satisfies the following two conditions:
@@ 2.1.1 Proximity oblivious testers
    If $f$ satisfies ${\mathcal{P}}$, then $\Pr\{\mathsf{T}^{f,D}=\text{Accept}\}=1$.
@@
    There exists a monotone function $\rho:(0,1]\to(0,1]$ such that if $f\not\in{\mathcal{P}}$, then $\Pr\{\mathsf{T}^{f,D}=\text{Reject}\}\geq\rho(\delta_{{\mathcal{P}},{\mathcal{D}}}(f))$.
The function $\rho$ represents the detection probability of the tester $\mathsf{T}$. Then we can obtain a one-sided error tester with parameter $\varepsilon$ by simply making $\Theta(1/\rho(\varepsilon))$ calls to the POT $\mathsf{T}$ and accepting if and only if all the calls return Accept (Goldreich and Ron, 2008).
For example, consider the BLR test (Blum et al., 1993) over the boolean hypercube ${\mathcal{X}}=\{0,1\}^{n}$, where the underlying distribution ${\mathcal{D}}$ is uniform. This is a one-sided error tester that uses $\varepsilon$ only to determine the number of calls it makes to a proximity oblivious tester. The POT $\mathsf{T}^{f,{\mathcal{D}}}$ is the machine that samples $X_{1},X_{2}\sim{\mathcal{D}}$ and then checks whether $f(X_{1})+f(X_{2})=f(X_{1}+X_{2})$ or not. Clearly if $f$ is indeed linear, then $\Pr\{\mathsf{T}^{f,D}=\text{Accept}\}=1$, while if $f$ is sufficiently far from ${\mathcal{P}}$, then there will be some nonzero probability of rejection by this tester (Goldreich, 2017).
@@ 2.2 Conformal prediction and conformal risk control
The conformal risk control framework extends conformal prediction to notions of error beyond miscoverage (Angelopoulos et al., 2024). Given an exchangeable collection of non-increasing, random loss functions $L_{i}:\Lambda\to(-\infty,B]$, $i=1,\ldots,n+1$, conformal risk control uses the first $n$ loss functions and calibration data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ to determine $\hat{\lambda}$ such that

### v2

@@ Abstract
AI model alignment is crucial due to inadvertent biases in training data and the underspecified machine learning pipeline, where models with excellent test metrics may not meet end-user requirements. While post-training alignment via human feedback shows promise, these methods are often limited to generative AI settings where humans can interpret and provide feedback on model outputs. In traditional non-generative settings with numerical or categorical outputs, detecting misalignment through single-sample outputs remains challenging, and enforcing alignment during training requires repeating costly training processes. In this paper we consider an alternative strategy. We propose interpreting model alignment through property testing, defining an aligned model $f$ as one belonging to a subset ${\mathcal{P}}$ of functions that exhibit specific desired behaviors. We focus on post-processing a pre-trained model $f$ to better align with ${\mathcal{P}}$ using conformal risk control. Specifically, we develop a general procedure for converting queries for testing a given property ${\mathcal{P}}$ to a collection of loss functions suitable for use in a conformal risk control algorithm. We prove a probabilistic guarantee that the resulting conformal interval around $f$ contains a function approximately satisfying ${\mathcal{P}}$. We exhibit applications of our methodology on a collection of supervised learning datasets for (shape-constrained) properties such as monotonicity and concavity. The general procedure is flexible and can be applied to a wide range of desired properties. Finally, we prove that pre-trained models will always require alignment techniques even as model sizes or training data increase, as long as the training data contains even small biases.
@@ 1 Introduction
The emergence of large foundation models has increased the attention to the problem of alignment. Aligned models are artificial intelligences designed to pursue goals that align with human values, principles, and intentions (Leike et al., 2018; Ouyang et al., 2022; Hendrycks et al., 2023; Ngo et al., 2024). Although the alignment problem is predominantly examined in the context of potential artificial general intelligence (AGI), large language models (LLMs), and reinforcement learning (RL) agents, it also has roots in the modern machine learning pipeline (D’Amour et al., 2022). Motivated by this, we introduce a broader notion of alignment in this paper, extending beyond the aforementioned generative models to include even tabular regression models.
@@
As an example, consider a regression task where property ${\mathcal{P}}$ represents models that are monotonically decreasing in a given feature (covariate). For example, predicting cancer patient survival should be monotonically decreasing in cancer stage (Vallon et al., 2022, 2024). Constraining a prediction model during training to maintain monotonicity in this feature can be viewed as a form of alignment. For a pre-trained model $f$ that was trained without such constraints, ensuring monotonically decreasing predictions in this feature can be more complex. This complexity arises particularly in non-generative settings where the user cannot update $f$ or obtain any outputs other than point predictions $f(X)$ for a given input $X$.
@@
In this work, we propose an approach to aligning a pre-trained model $f$ that is motivated by property testing (Ron, 2008; Goldreich, 2017) and conformal risk control (Angelopoulos et al., 2024). Property testing aims to design efficient algorithms for determining membership to the set ${\mathcal{P}}$ of functions with a given property, that require fewer resources than learning algorithms for ${\mathcal{P}}$ (Ron, 2008). This is particularly relevant for modern deep learning, where a user may need to determine if a pre-trained model $f$ belongs to ${\mathcal{P}}$ without the resources to train a model of comparable size.
@@
Property testing algorithms use local queries to determine, with high probability, whether a function has a given global property or is far from having it. We map such queries for a property ${\mathcal{P}}$ to a set of loss functions, which we then use in a conformal risk control procedure (Angelopoulos et al., 2024) to establish a notion of alignment for ${\mathcal{P}}$. We prove that this procedure yields a conformal interval around $f$ containing a function close to ${\mathcal{P}}$.
@@
We demonstrate our methodology on real-world datasets for the properties of monotonicity and concavity. Motivated by the potential for systematic under- or over-estimation bias in $f$, we provide a straightforward extension of Angelopoulos et al. (2024) to obtain asymmetric conformal intervals with multi-dimensional parameters. While we examine both monotonicity and concavity constraints, the majority of our focus is on monotonicity, as these constraints have been shown to promote crucial aspects of alignment to human values, such as fairness and adherence to social norms (Wang and Gupta, 2020).
@@
While our methodology provides a way to align pre-trained models, one may question whether such techniques will remain necessary as AI capabilities advance. Given the outstanding capabilities of modern AI models with substantially large numbers of parameters and training data, one may argue that the alignment problem may naturally disappear as such advances continue (Kaplan et al., 2020). However, another contribution of this paper is to refute this argument in a stylized setting, building on recent advances in the theory of linearized neural networks (Mei and Montanari, ; Misiakiewicz and Montanari, 2023). Specifically, we show that increasing the size of the training data or the number of parameters in a random feature model (a theoretically tractable neural network proxy where hidden layer weights are randomly initialized and fixed (Rahimi and Recht, 2007)) cannot help it satisfy a property ${\mathcal{P}}$, if the pre-training data has biased labels. Our simulations show that the result holds even if only a small fraction of the training labels are impacted by the bias.
@@
Summarizing our main contributions, we: (1) introduce an alignment perspective based on property testing, (2) use conformal risk control to post-process predictions of pre-trained models for better alignment, and (3) demonstrate that increasing training data and parameters in a random feature model does not eliminate the need for alignment. We discuss related work in Section 6, particularly our connections to Yadkori et al. (2024), who use conformal risk control to address large language model hallucinations (Ji et al., 2023).
@@ 2 Preliminaries
In this section we provide key definitions drawn from property testing as well as a condensed overview to conformal prediction and conformal risk control. We provide a short introduction to propery testing in Appendix A and an extensive introduction to the field can be found in Goldreich (2017).
2.1 Properties and Property Testing for Set-Valued Functions
Our perspective on alignment in this work is motivated by the field of property testing (Goldreich, 2017; Ron, 2008). Property testing studies algorithms that, by making a small number of local queries to a large object (such as a function or a graph), can determine whether the object has a certain property or is significantly far from having it.
Classic examples include linearity testing of Boolean functions (Blum et al., 1993), testing whether a function is a low-degree polynomial (Kaufman and Ron, 2006; Bhattacharyya et al., 2009), and testing $k$-juntas (Blais, 2009). These algorithms generally operate by randomly sampling and querying the object, leveraging local information to infer global properties.
In this work, we focus on set-valued functions, which are functions that map elements of a domain $\mathcal{X}$ to subsets of a codomain $\mathcal{Y}$, i.e., $F:\mathcal{X}\to 2^{\mathcal{Y}}$. While the standard definitions of property testing are technically sufficient for our purposes—since we can consider set-valued functions as functions with range $2^{\mathcal{Y}}$—we introduce specialized definitions to maintain clarity and to facilitate the transition between discussing $\mathcal{Y}$ and $2^{\mathcal{Y}}$.
Definition 1 (Satisfying and Accommodating a Property).
Let property $\mathcal{P}$ denote a specific subset of all functions that map $\mathcal{X}$ to $\mathcal{Y}$. A function $f:\mathcal{X}\to\mathcal{Y}$ satisfies the property $\mathcal{P}$ if $f\in\mathcal{P}$.
A set-valued function $F:\mathcal{X}\to 2^{\mathcal{Y}}$ accommodates a property $\mathcal{P}$ if there exists a function $g\in\mathcal{P}$ such that $g(x)\in F(x)$ for all $x\in\mathcal{X}$.
Intuitively, $F$ accommodates $\mathcal{P}$ if it contains at least one function $g$ satisfying $\mathcal{P}$ within its possible outputs.
We extend the notion of $\varepsilon$-farness from a property (as defined in Appendix A) to set-valued functions. For set-valued functions, we measure the distance based on how often the outputs of any function $g\in\mathcal{P}$ fall within the sets provided by $F$.
Definition 2 ($\varepsilon$-Faraway).
For a set-valued function $F:{\mathcal{X}}\to 2^{\mathcal{Y}}$, a distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, $\varepsilon>0$, and a property ${\mathcal{P}}$, we say $F$ is $\varepsilon$-Faraway from ${\mathcal{P}}$ with respect to ${\mathcal{D}}$ if $\delta_{{\mathcal{P}},{\mathcal{D}}}(F)>\varepsilon$, where
@@ 2.1 Properties and Property Testing for Set-Valued Functions
Note. Throughout this work, we assume that ${\mathcal{D}}$ is the empirical distribution of a fixed and finite calibration dataset, and thus has finite support. While this assumption is not strictly necessary, most property testing results are over finite domains. Property testing over functions with Euclidean domains is a in general a difficult problem, though there have been notable recent successes (Fleming and Yoshida, 2020; Arora et al., 2023)..
With these definitions in place, we can define testers for set-valued functions. We focus on one-sided error testers, which are algorithms that take in a set-valued function $F$, a distribution ${\mathcal{D}}$, and a distance parameter $\varepsilon$ and output either Accept or Reject. These algorithms never reject a function that accommodates the property. The standard definition of one-sided error testers (provided in Appendix A) extends naturally to set-valued functions by replacing the notion of satisfying a property with accommodating it.
Definition 3 (One-Sided Error Tester for Set-Valued Functions).
A one-sided error tester for a property $\mathcal{P}$ in the context of set-valued functions is a probabilistic oracle machine $\mathcal{M}$ that, given a distance parameter $\varepsilon>0$, oracle access to a set-valued function $F:\mathcal{X}\to 2^{\mathcal{Y}}$, and oracle access to samples from a fixed but unknown distribution $\mathcal{D}$ over $\mathcal{X}$, satisfies:
@@
    If $F$ accommodates $\mathcal{P}$, then $\Pr[\mathcal{M}^{F,\mathcal{D}}(\varepsilon)=\text{Accept}]=1$.
@@
    If $F$ is $\varepsilon$-Faraway from $\mathcal{P}$ with respect to $\mathcal{D}$, then $\Pr[\mathcal{M}^{F,\mathcal{D}}(\varepsilon)=\text{Accept}]\leq\frac{1}{3}$.
Here, $\mathcal{M}^{F,\mathcal{D}}(\varepsilon)$ denotes the execution of the tester $\mathcal{M}$ when given oracle access to the function $F$, the distribution $\mathcal{D}$, and the parameter $\varepsilon$.
Note that $\mathcal{M}$ itself is an abstract algorithm; $\mathcal{M}^{F,\mathcal{D}}$ is the instantiation of this algorithm with specific oracle access to $F$ and $\mathcal{D}$.
In many property testing algorithms, the parameter $\varepsilon$ is used only to determine the number of iterations or samples required, not the core logic of the tester. This leads to the concept of proximity-oblivious testers (POTs), where the basic testing procedure is independent of $\varepsilon$. The general definition of POTs (given in Appendix A) also extends naturally to set-valued functions.
Definition 4 (Proximity-Oblivious Tester for Set-Valued Functions).
A proximity-oblivious tester for a property $\mathcal{P}$ in the context of set-valued functions is a probabilistic oracle machine $\mathcal{T}$ that satisfies:
@@
    If $F$ accommodates $\mathcal{P}$, then $\Pr[\mathcal{T}^{F,\mathcal{D}}=\text{Accept}]=1$
@@
    There exists a non-decreasing function $\rho:(0,1]\to(0,1]$ (called the detection probability) such that if $F$ is $\varepsilon$-Faraway from $\mathcal{P}$,
    $\Pr[\mathcal{T}^{F,\mathcal{D}}=\text{Reject}]\geq\rho(\varepsilon).$
Here, $\mathcal{T}^{F,\mathcal{D}}$ denotes the execution of the tester $\mathcal{T}$ when given oracle access to the function $F$ and the distribution $\mathcal{D}$.
To obtain a one-sided error tester with parameter $\varepsilon$, we can make $\Theta\left(\frac{1}{\rho(\varepsilon)}\right)$ independent calls to the POT $\mathcal{T}$ and accept if and only if all the calls accept (Goldreich and Ron, 2008). We denote by $\mathcal{T}^{F,\mathcal{D}}(X)$ the output when applied to a specific sample $X\sim{\mathcal{D}}$, and note that with abuse of notation we will late consider ${\mathcal{D}}$ to be the empirical distribution of calibration dataset $\{(X_{i},Y_{i})\}_{i=1}^{n}$ in which case we write $\mathcal{T}^{F,\mathcal{D}}(X_{i},Y_{i})$ for the output on this specific sample from ${\mathcal{D}}$.
Example.
Consider functions $f:\mathbb{R}^{d}\to\mathbb{R}$, and let $\mathcal{P}$ denote the property that $f$ is constant in the $k$-th dimension. This property has has connections to fairness among other applications Caton and Haas (2024). Assume ${\mathcal{D}}$ is the empirical distribution of the inputs $X\in\mathbb{R}$ for some fixed dataset.
Restrict to set-valued functions $F$ that output compact and connected intervals of the form $[a,b]\subseteq\mathbb{R}$ for $a,b\in\mathbb{R}$. The candidate POT $\mathcal{T}^{F,\mathcal{D}}$ for whether such a set-valued function $F$ accommodates ${\mathcal{P}}$ is then as follows: sample $X,X^{\prime}\sim\mathcal{D}$, If $F(X)\cap F(X^{\prime})\neq\varnothing$, then Accept; otherwise, Reject. We prove that this satisfies Definition  4 in Appendix B.1.
@@ 2.2 Conformal prediction and conformal risk control
The conformal risk control framework extends conformal prediction to notions of error beyond miscoverage (Angelopoulos et al., 2024). Consider a paramater set $\Lambda\subset\mathbb{R}_{\geq 0}$ that is a bounded subset of the nonnegative reals. Given an exchangeable collection of non-increasing, random loss functions $L_{i}:\Lambda\to(-\infty,B]$, $i=1,\ldots,n+1$, conformal risk control uses the first $n$ loss functions and calibration data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ to determine $\hat{\lambda}$ such that